### v2

# Cyclicity of Multipliers on the Unit Ball of $\mathbb{C}^{n}$: A Corona-Based Approach

Abstract.
We study the cyclicity of multipliers in Dirichlet-type spaces $D_{\alpha}(\mathbb{B}_{n})$. Specifically, we show that a multiplier $f$ analytic on a neighborhood of the closed unit ball, whose zero set on the unit sphere is a compact, smooth, complex tangential submanifold of real dimension $m\leq n-1$, is cyclic in $D_{\alpha}(\mathbb{B}_{n})$ if and only if $\alpha\leq\frac{2n-m}{2}$, where $m$ is the real dimension of the zero set of $f$ on the boundary. Our approach combines classical results on peak sets in $A^{\infty}(\mathbb{B}_{n})$ due to Chaumat and Chollet with a Corona-type theorem with two generators for the multiplier algebra.
Key words and phrases:
Dirichlet-type Spaces, Weighted Bergman Spaces, Cyclic Vectors, Shift Operator
2020 Mathematics Subject Classification:
32F45
1. Introduction
Given a Banach space $\mathcal{A}$ of holomorphic functions in a domain in $\mathbb{C}^{n}$, suppose that polynomials are dense in $\mathcal{A}$ and act as multipliers - that is, for every polynomial $p$, the product $ph\in\mathcal{A}$ whenever $h\in\mathcal{A}$. A function $f\in\mathcal{A}$ is said to be cyclic if the linear span of all polynomial multiples of $f$ is dense in $\mathcal{A}$. Equivalently, $f$ is cyclic if there exists a sequence of polynomials $p_{n}$ such that $\|fp_{n}-1\|_{\mathcal{A}}\to 0$ as $n\to\infty$. The problem of characterizing cyclic functions is highly sensitive to the choice of space and remains nontrivial, even in classical Hilbert spaces of holomorphic functions on the unit disk. Complete classifications, such as Beurling’s theorem for Hardy space [5], are exceptional. The Brown–Shields conjecture, proposed in 1984 for the Dirichlet space, remains unresolved [2]. Recent progress in cyclicity in spaces of holomorphic functions on the unit disk, particularly Dirichlet-type spaces, can be found in [3], [9], [11], and [7]. A comprehensive textbook treatment of Dirichlet-type spaces is also available in [8].
In several variables, a complete characterization of cyclic functions in Dirichlet-type spaces remains elusive even for classical domains such as the unit ball and the polydisk. However, significant progress has been made in identifying cyclic polynomials, with explicit characterizations obtained in the unit ball of $\mathbb{C}^{2}$ and the bidisk [4], [13], [12]. Here, we prove a cyclicity result for multipliers whose zero set on the boundary has a specific structure. Our main result is the following:
Theorem 1.
Let $f\in\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))\cap\mathcal{O}(\overline{\mathbb{B}}_{n})$ be nonvanishing in the unit ball of $\mathbb{C}^{n}$, and suppose that $\mathcal{Z}(f)\cap\overline{\mathbb{B}}_{n}=M$, where $M$ is a compact complex tangential smooth manifold. Let $m\coloneqq\dim_{\mathbb{R}}(\mathcal{Z}(f)\cap\overline{\mathbb{B}}_{n})$, if $\alpha\leq\frac{2n-m}{2}$, then $f$ is cyclic in $D_{\alpha}(\mathbb{B}_{n})$.
A combination of the results in [17], which give a sufficient condition for non-cyclicity, and Theorem 1 leads to the following corollary.
Corollary 2.
Let $f$ be as in Theorem 1, then $f$ is cyclic in $D_{\alpha}(\mathbb{B}_{n})$ if and only if $\alpha\leq\frac{2n-m}{2}$.
Our main tool for approaching the problem is the following result. An analogous version in one variable was proved in [10] using Tolokonnikov’s Corona Theorem [16], and a slightly more general statement in several variables was recently obtained in [1]. We follow the same strategy here, using Theorem 8 to establish the result. This, together with the construction of a cyclic function $g$ in the multiplier space of $D_{\alpha}(\mathbb{B}_{n})$, denoted by $\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))$, dominated by $f$, will help us prove Theorem 1.
Theorem 3.
Let $f,g\in\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))\cap\mathcal{C}^{\infty}(\overline{\mathbb{B}}_{n})$ and assume that on $\mathbb{B}_{n}$ we have $|g|\leq|f|$ then there exists a natural number $N$ such that:
(1) $g^{N}]\subseteq[f]$
Here, by $[f]$ we mean the closed subspace of $D_{\alpha}(\mathbb{B}_{n})$ generated by $f$.
The organization of the paper is as follows. In Section 2, we develop the necessary tools to prove the main result. In particular, Subsection 2.1 introduces the notation and basic definitions; Subsection 2.2 is devoted to the proof of Theorem 3; and Subsection 2.3 presents several classical results of Chaumat and Chollet concerning peak sets of $A^{\infty}(\Omega)$, where $\Omega$ is a strictly pseudoconvex domain, and concludes with Corollary 10. These results are essential for constructing a cyclic function in the multiplier algebra with the desired zero set. In Section 3, we provide the proof of our main result, Theorem 1.
2. Preliminary Tools
2.1. Basics and Notations
Notation 1.
Let $\Omega\subseteq\mathbb{C}^{n}$ be a domain. We denote by $\mathcal{O}(\Omega)$ the set of all holomorphic functions $f:\Omega\to\mathbb{C}$.
Notation 2.
We write $A\lesssim B$ to indicate that there exists a constant $C>0$ such that $A\leq CB$. If both $A\lesssim B$ and $B\lesssim A$ hold, we write $A\asymp B$. The dependence of the constant $C$ on parameters will be specified whenever necessary.
Definition 1.
Let $\alpha\in\mathbb{R}$. A function $f\in\mathcal{O}(\mathbb{B}_{n})$, with power series expansion
$f(z)=\sum_{|k|=0}^{\infty}a_{k}z^{k},\quad z\in\mathbb{B}_{n},$
is said to belong to the Dirichlet-type space $D_{\alpha}(\mathbb{B}_{n})$ if
$\|f\|_{D_{\alpha}(\mathbb{B}_{n})}^{2}:=\sum_{|k|=0}^{\infty}(n+|k|)^{\alpha}\cdot\frac{(n-1)!}{k!\,(n-1+|k|)!}\,|a_{k}|^{2}<\infty,$
where $k$ is a multi-index, $|k|=k_{1}+\cdots+k_{n}$, $k!=k_{1}!\cdots k_{n}!$, and $z^{k}=z_{1}^{k_{1}}\cdots z_{n}^{k_{n}}$.
Definition 2.
Let $\alpha\in\mathbb{R}$. The multiplier space $\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))$ is the set of all functions $\varphi\in D_{\alpha}(\mathbb{B}_{n})$ such that for every $f\in D_{\alpha}(\mathbb{B}_{n})$, the product $\varphi f\in D_{\alpha}(\mathbb{B}_{n})$.
The space $\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))$ is a Banach algebra when equipped with the operator norm induced by multiplication. In particular, when $\alpha=0$, so that $D_{0}(\mathbb{B}_{n})=H^{2}(\mathbb{B}_{n})$, we have
$\mathcal{M}(D_{0}(\mathbb{B}_{n}))=H^{\infty}(\mathbb{B}_{n}),$
the space of bounded holomorphic functions on $\mathbb{B}_{n}$.
The following Lemma is a direct consequence of the fact that $D_{\alpha}(\mathbb{B}_{n})=A^{2}_{-(\alpha+1)}(\mathbb{B}_{n})$ in the sense of the generalization given to the definition of Weighted Bergman Spaces in [18].
Lemma 4.
$f\in D_{\alpha}(\mathbb{B}_{n})$ if and only if $R^{k}f\in D_{\alpha-2k}(\mathbb{B}_{n})$. Also whenever $\alpha<0$ we have:
(2) $f||^{2}_{\alpha}\simeq\int_{\mathbb{B}_{n}}(1-|z|^{2})^{-(\alpha+1)}|f(z)|^{2}\,dv(z)$
Corollary 5.
Let $\alpha\in\mathbb{R}$, and let $k\in\mathbb{N}$ be such that $\alpha-2k<0$. If $f\in D_{\alpha}(\mathbb{B}_{n})$ and $R^{k}f\in H^{\infty}(\mathbb{B}_{n})$, then $f\in\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))$.
Definition 3.
A function $f\in D_{\alpha}(\mathbb{B}_{n})$ is called a shift-cyclic vector if:
$f]:=\overline{\operatorname{span}\{pf:p\in\mathbb{C}[z_{1},\ldots,z_{n}]\}}=D_{\alpha}(\mathbb{B}_{n}).$
The next theorem provides a practical sufficient condition for the cyclicity of a function. For a detailed proof, one may consult [12].
Theorem 6.
Let $f\in D_{\alpha}(\mathbb{B}_{n})$ and assume that $f$ does not vanish on $\mathbb{B}_{n}$. For $0<r<1$, define the radial dilation
$f_{r}(z)=f(rz).$
If
$\sup_{0<r<1}\left\|\frac{f}{f_{r}}\right\|_{\alpha}<+\infty,$
then $f$ is cyclic in $D_{\alpha}(\mathbb{B}_{n})$.
2.2. Proof of Theorem 3
In order to prove Theorem 8 we need the following classical result on the solution of the $\bar{\partial}$-problem by Lieb and Range [15].
Theorem 7.
Let $\Omega\subset\subset\mathbb{C}^{n}$ be a strictly pseudoconvex domain with smooth boundary of class $\mathcal{C}^{k+2}$. Then for each $q=1,2,\ldots,n$, there exist linear operators
$T_{q}^{*}:\mathcal{C}_{(0,q)}^{\infty}(\overline{\Omega})\to\mathcal{C}_{(0,q-1)}^{\infty}(\overline{\Omega})$
with the following properties:
  1. (i)
    If $f\in\mathcal{C}_{(0,q)}^{\infty}(\overline{\Omega})$ and $\bar{\partial}f=0$, then
    $\bar{\partial}T_{q}^{*}f=f.$
  2. (ii)
    If $f\in\mathcal{C}_{(0,q)}^{k}(\overline{\Omega})$ with $\bar{\partial}f=0$, then for all $k=0,1,2,\ldots$,
    $\|T_{q}^{*}f\|_{\mathcal{C}^{k+1/2}(\overline{\Omega})}\leq C_{k}\|f\|_{\mathcal{C}^{k}(\overline{\Omega})}.$
The constants $C_{k}$ depend only on $k$ and $\Omega$, and are invariant under small $\mathcal{C}^{k+2}$-perturbations of $\Omega$.
Here, $\|\cdot\|_{\mathcal{C}^{k+1/2}(\overline{\Omega})}$ and $\|\cdot\|_{\mathcal{C}^{k}(\overline{\Omega})}$ denote the Hölder norms, where by $\|f\|_{k}$ and $\|f\|_{k+\varepsilon}$ we mean:
$\|f\|_{k}=\sum_{|\alpha|\leq k}\sup_{x\in\overline{\Omega}}|D^{\alpha}f(x)|,$
$\|f\|_{k+\varepsilon}=\|f\|_{k}+\sum_{|\alpha|=k}\sup_{x\neq y}\frac{|D^{\alpha}f(x)-D^{\alpha}f(y)|}{|x-y|^{\varepsilon}}.$
Theorem 8.
Let $f_{2},f_{1}\in\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))\cap\mathcal{C}^{\infty}(\overline{\mathbb{B}}_{n})$ and there exists $\delta>0$ such that $|f_{1}|^{2}+|f_{2}|^{2}\geq\delta$ on $\mathbb{B}_{n}$ then there exists $g_{1},g_{2}\in\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))\cap\mathcal{C}^{\infty}(\overline{\mathbb{B}}_{n})$ and natural number $k$ satisfying:
  1. (i)
    $f_{1}g_{1}+f_{2}g_{2}=1$.
  2. (ii)
    $\|g_{i}\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}\leq\frac{C}{\delta^{k}}$
for some constant $C>0$.
Proof.
Let $g_{i}^{0}=\frac{\bar{f}_{i}}{|f_{1}|^{2}+|f_{2}|^{2}}$ for $i=1,2$. Then $(g_{1}^{0},g_{2}^{0})$ satisfies (i). Note that
$\bar{\partial}g_{1}^{0}=f_{2}\omega,\quad\bar{\partial}g_{2}^{0}=-f_{1}\omega,$
where
$\omega=\frac{\bar{f}_{2}\,\bar{\partial}\bar{f}_{1}-\bar{f}_{1}\,\bar{\partial}\bar{f}_{2}}{(|f_{1}|^{2}+|f_{2}|^{2})^{2}}.$
Moreover, one easily observes that $\omega$ is a closed $(0,1)$-form, smooth up to the boundary, and all of its derivatives have $L^{\infty}$ norms bounded by constants of the form $\frac{C}{\delta^{k}}$, where $k$ depends on the order of the derivative. By Theorem 7, there exists a solution $\phi$ to the problem $\bar{\partial}\phi=\omega$, which is also smooth up to the boundary, and all of its derivatives have $L^{\infty}$ norms bounded by constants of the form $\frac{C}{\delta^{k}}$, again with $k$ depending on the differentiation order.
Now, we claim that the choice
$g_{1}=g_{1}^{0}-f_{2}\phi,\quad g_{2}=g_{2}^{0}+f_{1}\phi$
is the desired solution. Clearly, it satisfies (i). Moreover, by Lemma 4, a sufficient condition for $g_{1}$ and $g_{2}$ to be in the multiplier space of $D_{\alpha}(\mathbb{B}_{n})$ is that their radial derivatives up to a certain order are $L^{\infty}$-bounded. Hence, the norm estimate in (ii) follows. ∎
With the necessary tools in hand, we are now ready to achieve the main goal of this subsection:
Proof of Theorem 3.
For $\lambda\in\mathbb{C}$, define
$\delta(\lambda)=\inf_{z\in\mathbb{B}_{n}}\left\{|1-\lambda g(z)|^{2}+|f(z)|^{2}\right\}.$
If $|g(z)|\leq\frac{1}{2|\lambda|}$, then $|1-\lambda g(z)|^{2}\geq\frac{1}{4}$. On the other hand, whenever $|g(z)|\geq\frac{1}{2|\lambda|}$, we have $|f(z)|\geq\frac{1}{2|\lambda|}$. Therefore,
$\delta(\lambda)\geq\min\left\{\frac{1}{4|\lambda|^{2}},\frac{1}{4}\right\}.$
By Theorem 8, there exist $g_{1},g_{2}\in\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))$ such that:
  1. (A)
    $g_{1}(1-\lambda g)+g_{2}f\equiv 1$,
  2. (B)
    $\|g_{i}\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}\lesssim\frac{1}{\delta(\lambda)^{k}}\asymp|\lambda|^{2k}$ for $i=1,2$.
Now define the standard quotient map
$\pi:\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))\rightarrow\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))/[f]_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}.$
Then by (A),
$\pi(1-\lambda g)^{-1}=\pi(g_{1}).$
Thus, $\pi((1-\lambda g))^{-1}$, which is a holomorphic Banach-valued function of $\lambda$, is defined for every $\lambda$, so it is an entire function. Hence, we may write:
(3) $\|\pi((1-\lambda g))^{-1}\|=\left\|\sum_{j\geq 0}\lambda^{j}\pi(g^{j})\right\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))/[f]_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}}=\|\pi(g_{1})\|\leq\|g_{1}\|_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}\lesssim|\lambda|^{2k}.$
Now, by Liouville’s Theorem, we conclude that $\pi(g^{2k+1})=0$. In other words,
$g^{2k+1}]_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))}\subseteq[f]_{\mathcal{M}(D_{\alpha}(\mathbb{B}_{n}))},$
which implies
$g^{2k+1}]\subseteq[f].$
2.3. Peak Sets of $A^{\infty}(\mathbb{B}_{n})$
Definition 4.
Let $M\subset\partial\mathbb{B}_{n}$ be a smooth real submanifold of real dimension $m$. We say that $M$ is totally real if for every point $p\in M$, the complex structure $J$ satisfies
$T_{p}M\cap J(T_{p}M)=\{0\},$
where $J(v)=iv$ for $v\in\mathbb{C}^{n}$, and $T_{p}M$ is the real tangent space of $M$ at $p$.
Definition 5.
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded strictly pseudoconvex domain with smooth boundary $\partial\Omega$. A closed subset $E\subset\partial\Omega$ is called a peak set for the algebra $A^{\infty}(\Omega)$ if there exists a function $f\in A^{\infty}(\Omega)$ such that
$f(z)=1\quad\text{for all }z\in E,$
and
$|f(z)|<1\quad\text{for all }z\in\overline{\Omega}\setminus E.$
Equivalently, $E$ is a peak set if there exists a function $g\in A^{\infty}(\Omega)$ such that
$g(z)=0\quad\text{for all }z\in E,$
and
$\operatorname{Re}g(z)>0\quad\text{for all }z\in\overline{\Omega}\setminus E.$
Theorem 9 ([6], Theorem 21).
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded strictly pseudoconvex domain with $C^{\infty}$ boundary. Let $M\subset\partial\Omega$ be a $C^{\infty}$ submanifold of real dimension at most n1n-1 such that for every point $p\in M$, the tangent space $T_{p}(M)\subset T^{\mathbb{C}}_{p}(\partial\Omega)$. Then every compact subset $K\subset M$ is a peak set for the algebra $A^{\infty}(\Omega)$.
Moreover, if $g$ is the corresponding peak function in the sense of Definition 5, then there exists a constant $c>0$ and an open neighborhood $U\subset\mathbb{C}^{n}$ of $K$ such that
(4) $\operatorname{Re}g(z)\geq c\cdot\operatorname{dist}(z,M)^{2}\quad\text{for all }z\in\Omega\cap U.$
Corollary 10.
Under the hypotheses of Theorem 9, for every point $z_{0}\in K$, there exist constants $c>0$ and $c^{\prime}>0$, and an open neighborhood $U\subset\mathbb{C}^{n}$ of $z_{0}$, such that for all $z\in U\cap\Omega$, we have
$|g(z)|\geq c\left(u+|v|+\mathrm{dist}(z,M)^{2}\right),$
and
$|g(z)|\leq c^{\prime}\,\mathrm{dist}(z,M),$
where $w=u+iv$ are local holomorphic coordinates centered at $z_{0}$, with $u$ in the inward normal direction to $\partial\Omega$, and $v$ in the tangential direction.
Proof.
Note that $\operatorname{Re}g$ is a non-constant pluriharmonic function that attains its absolute minimum on the boundary point $z_{0}$. By Hopf’s Lemma, we have
$\left.\frac{\partial}{\partial u}\operatorname{Re}g\right|_{z_{0}}\neq 0.$
Applying the Cauchy–Riemann equations then yields
$\left.\frac{\partial}{\partial v}\operatorname{Im}g\right|_{z_{0}}=\left.\frac{\partial}{\partial u}\operatorname{Re}g\right|_{z_{0}}\neq 0.$
This, together with (4), completes the proof. ∎
3. Proof of the Main Result
Lemma 11.
Let $g\in A^{\infty}(\mathbb{B}_{n})$ be non-vanishing on $\mathbb{B}_{n}$ then for every positive integer $l$ there exists a constant $C_{l}>0$ independent of $r$ such that we have:
(5) $|R^{l}\frac{g}{g_{r}}|\leq\frac{C_{l}(1-r)}{|g_{r}|^{l+1}}$
Proof.
Let $g=\sum_{k\geq 0}g_{k}$ where $g_{k}$’s are $k$-homogeneous polynomials. So if we define $\phi=g_{r}Rg-gRg_{r}$, we may write:
$|R^{t}\phi|=|R^{t}((g_{r}-g)Rg+g(Rg-Rg_{r}))|=|R^{t}(Rg\sum_{k\geq 0}(r^{k}-1)g_{k}+g\sum_{k\geq 0}(1-r^{k})kg_{k})|\leq C_{t}(1-r)$
where $C_{t}$ depends only on the order of differentiation $t$. Noting that $|R^{l}\dfrac{g}{g_{r}}|=|R^{l-1}\dfrac{\phi}{g_{r}^{2}}|$ finishes the proof of the lemma since $g$ and all of its derivatives are bounded. ∎
Lemma 12.
Let $M\subset\partial\mathbb{B}_{n}$ be as in the Theorem 9 with $m\coloneqq\dim_{\mathbb{R}}M\leq n-1$ and $K\subseteq M$ compact then there exists a function $g\in A^{\infty}(\mathbb{B}_{n})$ such that $\mathcal{Z}(g)\cap\overline{\mathbb{B}}_{n}=K$ which is cyclic in $D_{\alpha}(\mathbb{B}_{n})$ for any $\alpha\leq\alpha_{c}=\frac{2n-m}{2}$. In fact, the function $g$ here is the same as the one in Theorem 9.
Proof.
Corollary 10 ensures the existence of a function $g\in A^{\infty}(\mathbb{B}_{n})$ with the desired zero set. We will show that this function is cyclic in $D_{\alpha_{c}}(\mathbb{B}_{n})$.
Let $r\in(0,1)$, and define $g_{r}(z)=g(rz)$. Our goal is to prove that
$\sup_{r\in(0,1)}\left\|\frac{g}{g_{r}}\right\|_{\alpha_{c}}<\infty,$
so that we can invoke Theorem 6 to conclude the proof.
Let $l$ be a sufficiently large positive integer such that $2l-\alpha_{c}>0$. Then, by Lemma 4, we have:
(6) $\left\|\frac{g}{g_{r}}\right\|_{\alpha_{c}}\asymp\int_{\mathbb{B}_{n}}(1-|z|^{2})^{2l-\alpha_{c}-1}\left|R^{l}\left(\frac{g}{g_{r}}\right)\right|^{2}\,dv(z).$
Applying Lemma 11 to (6), we obtain:
(7) $\left\|\frac{g}{g_{r}}\right\|_{\alpha_{c}}\lesssim(1-r)^{2}\int_{\mathbb{B}_{n}}(1-|z|^{2})^{2l-\alpha_{c}-1}\left|\frac{1}{g_{r}}\right|^{2l+2}\,dv(z).$
Now, for $p\in M$, let $Z(x,y,w)$ be a local coordinate at $p$, with $x\in\mathbb{R}^{m}$, $y\in\mathbb{R}^{2n-m-2}$, and $w=u+iv\in\mathbb{C}$, such that $Z(x,0,0)\in M$, $Z(x,y,0)\in\partial\mathbb{B}_{n}$, and $w$ is as in Corollary 10. Then, for $r>\frac{1}{2}$, there exists $c>0$ such that
(8) $|g_{r}(Z)|\geq c\left((1-r)+u+|v|+|y|^{2}\right).$
To estimate the integral in (6), it suffices to restrict it to the region where the local coordinate $Z$ is valid. Thus, we write:
(9) $\left\|\frac{g}{g_{r}}\right\|_{\alpha_{c}}\lesssim(1-r)^{2}\int_{(0,1)^{m}}\int_{(0,1)^{2n-m-2}}\int_{(-1,1)}\int_{(0,1)}\frac{u^{2l-\alpha_{c}-1}}{((1-r)+u+|v|+|y|^{2})^{2l+2}}\,du\,dv\,dy\,dx.$
Letting $u=u^{\gamma}$ where $\gamma=2l-\alpha_{c}$, we may rewrite (9) as
$\left\|\frac{g}{g_{r}}\right\|_{\alpha_{c}}$ $\lesssim(1-r)^{2}\int_{(0,1)^{m}}\int_{(-1,1)^{2n-m-2}}\int_{(-1,1)}\int_{(0,1)}\frac{1}{((1-r)+u^{1/\gamma}+|v|+|y|^{2})^{2l+2}}\,du\,dv\,dy\,dx$
$\asymp(1-r)^{2}\int_{(0,1)^{m}}\int_{(-1,1)^{2n-m-2}}\int_{(-1,1)}\int_{(0,1)}\frac{1}{((1-r)^{\gamma}+u+|v|^{\gamma}+|y|^{2\gamma})^{\frac{2l+2}{\gamma}}}\,du\,dv\,dy\,dx$
$\asymp(1-r)^{2}\int_{(0,1)^{m}}\int_{(-1,1)^{2n-m-2}}\int_{(-1,1)}\frac{1}{((1-r)^{\gamma}+|v|^{\gamma}+|y|^{2\gamma})^{\frac{2l+2}{\gamma}-1}}\,dv\,dy\,dx$
$\asymp(1-r)^{2}\int_{(0,1)^{m}}\int_{(-1,1)^{2n-m-2}}\int_{(-1,1)}\frac{1}{((1-r)+|v|+|y|^{2})^{\alpha_{c}+2}}\,dv\,dy\,dx$
$\asymp(1-r)^{2}\int_{(0,1)^{m}}\int_{(-1,1)^{2n-m-2}}\int_{(0,1)}\frac{1}{((1-r)+v+|y|^{2})^{\alpha_{c}+2}}\,dv\,dy\,dx$
$\asymp(1-r)^{2}\int_{(0,1)^{m}}\int_{(-1,1)^{2n-m-2}}\frac{1}{((1-r)+|y|^{2})^{\alpha_{c}+1}}\,dy\,dx$
$\asymp(1-r)^{2}\int_{(0,1)^{m}}\int_{(-1,1)^{2n-m-2}}\frac{1}{((1-r)+y_{1}^{2}+\cdots+y_{2n-m-2}^{2})^{\alpha_{c}+1}}\,dy\,dx$
$\asymp(1-r)^{2}\int_{(0,1)^{m}}\int_{(0,1)^{2n-m-2}}\frac{1}{((1-r)^{1/2}+y_{1}+\cdots+y_{2n-m-2})^{2n-m+2}}\,dy\,dx$
$\asymp(1-r)^{2}\int_{(0,1)^{m}}\frac{1}{(1-r)^{2}}\,dx$
$=1.$
Proof of Theorem 1.
By Theorem 9, we know that $\mathcal{Z}(f)\cap\overline{\mathbb{B}}_{n}=M$, is a peak set by Theorem 9. Therefore, by Lemma 12, there exists $g\in A^{\infty}(\mathbb{B}_{n})$ such that $g$ is cyclic in $D_{\frac{2n-m}{2}}(\mathbb{B}_{n})$, $\mathcal{Z}(g)\cap\overline{\mathbb{B}}_{n}=M$, and $g$ is a cyclic multiplier.
Note that for $g$, by construction, we have:
$|g(z)|\lesssim\operatorname{dist}(z,M).$
On the other hand, Łojasiewicz’s inequality [14] gives that there exists a positive integer $l$ and a constant $C>0$ such that for all $z\in\overline{\mathbb{B}}_{n}$ we have:
$|f(z)|\geq C\operatorname{dist}(z,M)^{l}.$
Therefore, possibly after raising $g$ to a suitable power and multiplying by a positive constant, we have on $\mathbb{B}_{n}$:
$|g(z)|\leq|f(z)|.$
Noting that $f$ and $g$ satisfy the conditions of Theorem 3, and that $g$ is cyclic (hence so is $g^{N}$, since $g$ is a multiplier), completes the proof. ∎
Acknowledgements
The author would like to express sincere gratitude to Professors Łukasz Kosiński and Pascal Thomas, the author’s PhD advisors, for their invaluable guidance and ongoing support throughout this work. The author is particularly grateful to Professor Thomas for his substantial help in the development of the present idea, and to Professor Kosiński for his continued insightful advice and encouragement. Finally, the author acknowledges the Institut de Mathématiques de Toulouse and the Department of Mathematics at Jagiellonian University for providing a stimulating research environment.